OpenStudy (anonymous):

The shipping company BigBlue has a size requirement for rectangular packages: the sum of the three dimensions cannot exceed 135 inches. If one end of the package is a square with sides of x inches, and the other dimension is y inches, what are the dimensions that will maximize the volume?

OpenStudy (anonymous):

Is this calculus? My main question is can I solve it using calculus?

OpenStudy (anonymous):

yes, i think the first derivative too. but if you cant just tell me what you know

OpenStudy (anonymous):

I can. I think I got it. Are you ready?

OpenStudy (anonymous):

yayy

OpenStudy (anonymous):

we know that 2x + y = 135

OpenStudy (anonymous):

Vmax = (x^2)y

OpenStudy (anonymous):

Lets get the volume equation in terms of x

OpenStudy (anonymous):

so we solve the first and substitute. Vmax=(x^2)(135-2x)

OpenStudy (anonymous):

y=135-2x?

OpenStudy (anonymous):

lets distribute x^2... Vmax=135x^2-2x^3

OpenStudy (anonymous):

Are you with me so far? or is there is step that confuses you?

OpenStudy (anonymous):

x = 45 and 0

OpenStudy (anonymous):

no go on

OpenStudy (anonymous):

not yet. I think we must take the derivative next and find x when the slope = 0

OpenStudy (anonymous):

1st derivative is 270x-6( x^2)

OpenStudy (anonymous):

so Vmax' =270x-6x^2

OpenStudy (anonymous):

nice.

OpenStudy (anonymous):

Set Vmax' = 0 and solve for x

OpenStudy (anonymous):

yep x= 45 and 0

OpenStudy (anonymous):

ok. haha nice. I didn't know you were that far already.

OpenStudy (anonymous):

i just needed to go past the substitution, thank you... Gosh ill love you forever

OpenStudy (anonymous):

i guess x=0 is extraneous

OpenStudy (anonymous):

Your welcome. I guess the final anser would be 45 x 45 x 45

OpenStudy (anonymous):

y= 135-2(45)

OpenStudy (anonymous):

I wouldn't have guessed a perfect cube in the beginning but I think the math is all right.

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

how do i find the The maximal volume ?

OpenStudy (anonymous):

45*45*45?

OpenStudy (anonymous):

Well we just found the x value where the volume would be at its maximum. So now just use V=lwh. So 45^3

OpenStudy (anonymous):

can i throw in another question?

OpenStudy (anonymous):

its alittle tricky and i got stuck?

OpenStudy (anonymous):

A crate open at the top has vertical sides, a square bottom, and a volume of 864 cubic meters. If the crate has the least possible surface area, find its dimensions. Hint : Each surface of the box is what shape?

OpenStudy (anonymous):

first thing we need to take into account is there is no top

OpenStudy (anonymous):

V=864cm^3 V=hx^2 SA= x^2 + 4xh

OpenStudy (anonymous):

Those are the equations I came up with based on the information given

OpenStudy (anonymous):

i always have a problem figuring out how to do the substitution

OpenStudy (anonymous):

ok. So solve the Volume equation for h and tell me what you get

OpenStudy (anonymous):

so i guess h = 864cm^2/x^2

OpenStudy (anonymous):

sorry cm^3

OpenStudy (anonymous):

ya so that IS h. so rewrite the SA equation and where there is an h put what it is in terms of x

OpenStudy (anonymous):

do the powers subtract? since eventually x^2 is cm^2 ?

OpenStudy (anonymous):

leave off the units if it is confusing you for now

OpenStudy (anonymous):

SA(x) = x^2+ 3456x/x^2

OpenStudy (anonymous):

ok. Do you see anywhere where that equation can be simplified

OpenStudy (anonymous):

2x-3456x^-1 ?

OpenStudy (anonymous):

S = x^2 + 3456x^-1 S' = 2x - 3456x^-2 S' = 2x - 3456/x^2

OpenStudy (anonymous):

3456=2x^3 ?

OpenStudy (anonymous):

Thats what I did

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

Plug it into our original equations and I think we have our answer

OpenStudy (anonymous):

y=6

OpenStudy (anonymous):

Correcttttt!!!!

OpenStudy (anonymous):

I have to head to class. Good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

good luck to you too